Kelvin W. Sherlock

                                                    ksherlock@gmail.com

Education University of Vermont, Burlington, Vermont

■ 1995–1999 B.S. in Biomedical Technology Minor in Computer Science GPA 3.62

■ 2000 Business Law I Business Law II GPA 4.0

■ 2001 Quality Management GPA 4.0

■ 2003 Compiler Construction GPA 4.0

Champlain College, Burlington, Vermont ■ 2002

Microeconomics GPA 4.0

University of Florida, Gainesville, Florida ■ 2008

Database Management Systems GPA 4.0

Epic Systems Corporation, Verona, Wisconsin ■ Certified in:

Bridges Interfaces (2008, 2009, 2010,2012) Cadence (2009, 2010, 2012) Chronicles Install Utilities (2009, 2010, 2012) Chronicles Non-Programming (2009, 2010, 2012) Chronicles Programming (2009, 2010, 2012) Clarity ETL Administration (2008, 2009, 2010, 2012) Data Courier Administrator (2009, 2010, 2012) Hyperspace Configuration (2009, 2010, 2012) Prelude (2009, 2010, 2012) Welcome (2010, 2012)

Work Experience

Application Analyst L2  University of Florida Physicians
Gainesville, Florida  April 2007–present

• Analyzed applications.

• Provided support and AES programming for Flowcast/Centricity applications (BAR, ETM, HCO, IDXWeb, PCS, Sched, SIU and TES).

• Designed and implemented web-based applications with ASP, ASP.NET, and SQL Server.

• Lead Epic Cadence and Prelude implementations.

• Assisted with Epic Ambulatory EMR implementation.

• Assisted with Epic Resolute Implementation.

• Supported other Epic applications, including work with Bridges, Clarity, Data Courier, Identity and creating custom programming points.

• Converted IDX appointments, guarantors, and coverages to Epic.

• Created custom Caché and VB applications for viewing and editing Epic Chronicles Data.

Contract Programmer  Henry Elliott/Coventry Health Care
Cranberry, Pennsylvania  August 2006

• Assisted with project for installation and customization of Medicaid HMO software.

• Provided technical information to other employees.

Web Developer  2005–present

• Designed, built, and maintained web sites with XHTML, Flash, JavaScript, PHP, perl, DHTML, AJAX, and SQL (MySQL, PostgreSQL).

• Provided maintenance, support, and customizations for e-commerce applications.

Customer Support Software Engineer  IDX Systems Corporation
South Burlington, Vermont  September 1999–October 2004

• Provided functional and technical support for the M-based IDXtend/Flowcast/Centricity suite of applications (including BAR, EWS/SCHED, SIU, TES, MCA, and VM/HPA).

• Designed and implemented custom enhancements to SCHED, EPMS, and VM.

• Wrote conversion utilities used to import registration and insurance information from legacy systems into Flowcast.

• Wrote conversion utilities used to export registration, insurance, BAR charges, and HPA charges for use with other systems.

• Wrote, installed and supported HL7 interfaces used to send and receive ADT, charge, dictionary, insurance, registration, and scheduling information.

• Designed and implemented custom interfaces used to send charges to bad debt agencies.

• Provided client consultation for HL7 performance and modifications.

• Repaired low-level and high-level database and integrity errors.

• Recreated failures and proposed fixes for future releases.

Information Counselor  Computer Information Technology, UVM
Burlington, Vermont  September 1995–June 1999

• Answered client questions regarding Macintosh, Unix, and Windows applications.

• Administered a Linux server.

Software Developer  1996–present

• Designed and implemented custom programs and utilities for iPhone, Macintosh, Linux, FreeBSD, BeOS, Windows, and Apple IIgs with C, C++, Objective C, and Assembly language.

• Developed games and utilities for the iPhone, iPod Touch, and iPad.

• Developed Silver Platter, a commercially available HTTP server.

• Developed BShisen, a popular BeOS game which was named the BeWare Gem of the Week by Be Inc.

• Developed GShisen, a popular Apple IIgs game which received the 1998 Apple II Achievement Award for Best Freeware Product.

Publications Sherlock, Kelvin. “Window Menu Bars.” Juiced.GS 12.1 (2007).

Sherlock, Kelvin. “ProFUSE: A ProDOS Filesystem in Userspace.” Juiced.GS 14.1 (2009).

Skills Programming Languages: ActionScript, ASP/ASP.NET C, C++, C♯, Objective C, Erlang, Java, JavaScript/JScript, Pascal, Perl, PHP Markup Languages: CSS, HTML, TEX/L^{A}TEX, XML, XSL Assembly Languages: 6502, 65816, x86, ARM Operating Systems: BSD, Linux, Macintosh OS X, Solaris, VMS, Windows NT Databases: M/MUMPS (Caché, ISM, DSM), MySQL, PostgreSQL, SQL Server